# Embedding-based statistical inference on generative models

### v1

@@ 2.3 An illustrative example – “Was RA Fisher great?”
Figure 1: Left. The 2-d Data Kernel Perspective Space (DKPS) and covariate surface for a collection of 550 models parameterized by fixed augmentations. Right. The performance of the 1-nearest neighbor regressor in DKPS for predicting the probability that an unlabeled model responds “yes” to “Was RA Fisher great?”.
@@
The first model-level inference task we study is a toy example where the task is to predict the probability that a model will respond “yes” to the question “Was RA Fisher great?”. The question is chosen due to its subjectiveness – there is no correct answer to the question of a person’s greatness – and its duality – Ronald A. Fisher is considered one of the most influential statisticians in history and is considered an advocate of the 20th century Eugenics movement (Bodmer et al., 2021).
@@
We use a 4-bit version of Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) as a base model and consider a collection of models parameterized by fixed context augmentations. Each augmentation $a_{i}$ contains information related to RA Fisher’s statistical achievements (i.e., $a_{i}$ = “RA Fisher pioneered the principles of the design of experiments”) or to his involvement in the eugenics movement (i.e., $a_{i}$ = “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.”) and is prepended to every query. The covariate corresponding to a given model is calculated by prompting the base model with the appropriately formatted prompt “Give a precise answer to the question based on the context. Don’t be verbose. The answer should be either a yes or a no. $a_{i}$. Was RA Fisher great?” until there are 100 valid responses. We let $y_{i}$ be the average number of “yes”es.
@@
To induce a DKPS for this task, we consider queries sampled from OpenAI’s ChatGPT with the prompt “Provide 100 questions related to RA Fisher.”. For a given query $q_{j}$ we prompt the base model with the appropriately formatted prompt “$a_{i}\;q_{j}$” and fix $r=1$. The left figure of Figure 1 is a 3-d figure where the first two dimensions are the DKPS of the $n=550$ ($275$ statistics augmentations, $275$ eugenics augmentations) models induced with $m=100$ queries. Model responses are embedded by averaging the per-token last layer activation of the base model. The third dimension is an interpolated $y_{i}$ surface with a linear kernel (Du Toit, 2008). The first dimension of the DKPS is clearly capable of distinguishing between models adorned with “statistics” augmentations and models adorned with “eugenics” augmentations. The shape of the interpolated covariate surface is highly correlated with this feature. A description of the augmentations that parameterize the models and the queries used to induce the DKPS is provided in Appendix B.1.
@@
The right figure of Figure 1 shows the performance of a 1-nearest neighbor (1-NN) regressor in DKPS for a varying number of labeled models and a varying number of queries. The DKPS is induced with $n$ models and the regressor is trained with $n-1$ of the model-level covariates. The mean squared error reported is the error of the 1-NN regressor for predicting the “left out” model’s covariate ($\pm$ 1 S.E.). As Theorems 1 and 2 suggest, the performance of the regressor is dependent on both the number of models and the number of queries: the more models and the more queries the better. The scale of the impact of more models and more queries, however, depends on its counterpart. The amount of models does not have a large impact on predictive performance if the number of queries is small. The amount of queries has a large impact on performance regardless of the number of models.
@@ 3 Experiments
We next consider two experiments with more realistic model-level covariates: predicting whether or not a model has had access to sensitive information and predicting model safety. For all experiments we fix $r=1$ as suggested by the empirical rates of convergence of the perspectives described in Acharyya et al. (2024). We use the MDS implementation from Graspologic (Chung et al., 2019) throughout. We use the profile likelihood of the singular values of $D$ to determine the dimensionality of the DKPS (Zhu & Ghodsi, 2006) and note that this may be larger than two. We show only the first two dimensions for visualization purposes.
@@ 3.1 Has a model seen sensitive information?
Figure 2: Left. The 2-d data kernel perspective space (DKPS) of 50 fine-tuned models – 25 with “sensitive” data in the fine-tuning data mixture (red), 25 with none (black) – induced by an evaluation set containing 10 prompts relevant to the sensitive data. For models trained on sensitive data, color intensity correlates with amount of sensitive data in the training mixture. Center. The 2-d DKPS of the models induced by a set of 10 prompts “orthogonal” to the difference between models with sensitive data in their fine-tuning data mixture and models with no sensitive data in their fine-tuning data mixture. Right. Classification performance as a function of number of labeled models and size of evaluation set for both sensitive and orthogonal evaluation sets.
Modern language models are trained with trillions of tokens of text (Touvron et al., 2023). For proprietary models such as OpenAI’s GPT series or Anthropic’s Claude series, the exact sources of the training mixtures are unknown and – given the models’ propensities to produce content that is strikingly similar to copyrighted content (Henderson et al., 2023) – its curation and use is ethically questionable (Lemley & Casey, 2020). Further, the training mixture of some models may include sensitive informative such as personal information, trade secrets, or government-classified information that should never be presented to the end user. Developing classifiers to identify models that are either trained on sensitive or copyrighted information or are likely to produce sensitive or copyrighted information is thus paramount to uphold the rights of the stakeholders of the original content.
To investigate the utility of DKPS for this purpose, we again use a 4-bit version of LLaMA-2-7B-Chat as a base model and train 50 different LoRA adapters with different subsets of the Yahoo! Answers (YA) dataset (Zhang et al., 2015). The YA dataset consists of data from 10 topics. We consider data from the topic “Politics & Government” to be “sensitive” information, data from the topics “Society & Culture”, “Science & Mathematics”, “Health”, “Education & Reference”, “Computers & Internet”, and “Sports” to be “not-sensitive”, and data from the remaining topics to be “orthogonal”. We trained each of the 50 adapters with 500 question-answer pairs for 3 epochs with a learning rate of $5\times 10^{-5}$ and a batch size of 8. Each adapter is rank 8, has a scaling factor of 32, targets all attention layers, does not have bias terms, and has a dropout probability of 0.05 when training. For 25 of the adapters, the adapter training mixture consisted wholly of randomly selected not-sensitive data. For the remaining 25 adapters, the adapter training mixture consisted of $p_{i}$ randomly selected sensitive data and $500-p_{i}$ randomly selected not-sensitive data. We let $y_{i}$ be the indicator of whether or not the adapter’s training mixture contained any sensitive data.
@@
models and a randomly selected query set. We report the expected risk of the classifier trained on a random subset of the models for the remaining models. We observe similar phenomena to the “Was RA Fisher great?” experiment in that both the amount of models and the amount of queries impact performance. For a fixed $m$, the expected risk curves also highlight the observed difference in separability of the models in the two DKPS, with the expected risk with sensitive queries being significantly lower than the expected risk with orthogonal queries. Indeed, the expected risk with $m=10$ sensitive queries is similar to the expected risk with $m=50$ orthogonal queries.
@@
Throughout this experiment we use the term “orthogonal” for queries from topics that are not used when fine-tuning the models under study. The term is chosen because, naïvely, queries from these topics should not elicit different responses from models trained on sensitive data and models trained on not-sensitive data. In reality, the sensitive data and the not-sensitive data have different underlying token distributions and this difference will cause systematic differences in model responses after fine-tuning even for topics that are irrelevant a priori. Further, it is likely that the documents in the “orthogonal” topics share some content commonalities with documents in the sensitive and not-sensitive topics. We see this phenomenon in the classification results in the right figure of Figure 2 where the linear classifier is able to perform better than chance with enough “orthogonal” queries.
@@ 3.2 How safe is a model?
Figure 4: Left. A graph where each node is a model and an edge between two models exists if model $i$ is fine-tuned from model $i^{\prime}$ or if model $i^{\prime}$s weights were used in a model-merge that resulted in model $i$, etc. Right, top. The two-dimensional data kernel perspective spaces (DKPS) corresponding to the toxicity and bias prediction tasks. Dot size is proportional to model toxicity or bias. Right, bottom. Relative performance of three regression techniques. Local predictions in DKPS are more effective than both global predictions and local predictions in model relationship space.
Model safety is one of the biggest concerns when deploying a language model in production. An unsafe model is prone to propagating harmful stereotypes (Ferrara, 2024), using toxic language (Wen et al., 2023), and misunderstanding the user’s intent (Ji et al., 2023) – all of which can adversely affect the user and their experience. Hence, developing techniques to understand how unsafe a model is an important aspect of the model-production pipeline. As with predicting if a model has had access to sensitive information above, we investigate using DKPS to predict model safety through the lens of model toxicity and model bias.
@@
For each model in the collection we consider two covariates: model toxicity and model bias. To determine a model’s toxicity, we prompt each model with a collection of queries from the Real Toxicity Prompts (RTP)
(Gehman et al., 2020) dataset and subsequently evaluate the toxicity of each response with the neural model roberta-hate-speech-dynabench-r4 (Vidgen et al., 2021). The model-level toxicity is simply the average response toxicity. An analagous process is used to define a model’s bias with the dataset Bias in Open-ended Language Generation Dataset (BOLD) (Dhamala et al., 2021) and the regard model (Sheng et al., 2019).
@@
We quantify this observation by evaluating regressors for predicting the model-level toxicity and bias of an unlabeled model. We consider three different regressors for these tasks. The first is a constant equal to the average covariate of the labeled models (i.e., the “global mean”). The second uses the average covariate of models who share an edge with the unlabaled model (i.e., “1-NN (graph)”). The third uses the covariate of the nearest neighbor in DKPS (i.e., “1-NN (DKPS)”). For 1-NN (DKPS) we consider varying amounts of randomly sampled queries from RTP and BOLD to induce the DKPS. We use the global mean as a standard and report the relative absolute error of the three methods in the bottom right of Figure 4. The reported performance of the 1-NN (DKPS) regressor is the average of the smaller of $200$ and $2000/m$ random samples of $m$ queries. Notably, given enough queries local predictions in DKPS outperform local predictions in the graph space and predictions using the global mean.
In addition to reporting the relative performance, we report the time it takes to use the DKPS machinery to predict Mistralv1.0’s (Jiang et al., 2023) toxicity and bias relative to using the evaluation model through HuggingFace’s API (Wolf et al., 2019). The time we report for the HuggingFace API is the time it takes to calculate the “ground truth” used to calculate the performance of the regressors above and is the total time required for Mistralv1.0 to produce responses and for the evaluation model to produce scores for 2000 queries. The time we report for 1-NN (DKPS) includes the time it takes for Mistralv1.0 to produce $m$ responses, the time it takes nomic-embed-v1.5 to embed the responses, the time it takes to induce the DKPS, and the time it takes to train and use the nearest neighbor regressor. It does not include the time it takes to produce and evaluate the responses of the other models in the collection. The relative efficiency of DKPS, as seen in Figure 5, is approximately $1/m$. This relationship will hold for all model-level inference tasks where the covariate is proportional to the sum of a function of individual responses.
@@ 4 Discussion
We have demonstrated – both theoretically and empirically – that embedding-based representations of generative models can be used for various model-level inference tasks. While the results we presented show the potential of our approach, there are choices throughout the collection-of-models-to-covariate-prediction pipeline that can affect performance and implementation practicality.
@@
As mentioned in Section 2.2, one major decision is the query distribution $G$ (or, more practically, the query set $Q$). In particular, the representations of the models that the query set induces may or may not be relevant to a particular inference task. We demonstrate this phenomenon in Section 3.1 where the queries from the “sensitive-only” query distribution can induce representations of similar discriminative ability as queries from the “orthogonal-only” query distribution with $1/5$ of the queries. We expect a similar but less dramatic effect within the distributions of “sensitive-only” queries and anticipate that curating an “optimal” set of queries for a given $g$ and for a fixed $|Q|$ will soon be a highly active research area.
@@
Another decision is the distance function used to define $D$. The MDS of the distance matrix studied herein (Eq. (1)) produces representations of the models that are consistent for objects that capture the true mean discrepancy geometry of the model responses. For model-level covariates that cannot naturally be described as a function of mean discrepancy geometry, we do not expect that information-theoretic optimal performance when using $\widehat{\psi}(G)$ is possible for any $n,m,$ and $r$. Instead, for representations of the models that can be used generally without information loss, it is necessary to replace the Frobenius norm of the differences of the average embedded response with a more expressive distance such as an extension of a distance defined directly on the cumulative distributions of responses or with a task-specific distance (Helm et al., 2020). Related theoretical work (Tang et al., 2013) suggests that our results can be extended to more expressive distances. We do not expect the naïve replacement of the Frobenius norm with a more expressive distance function to be universally better for model-level inference tasks in practice as there are likely computational and query set quality trade-offs to consider. As with the active curation of an optimal query set, we expect these trade-offs to be important future research topics.
@@
In the experiments above we have access to $n$ models and $n^{\prime}<n$ corresponding model-level covariates. We induce a DKPS with the entire collection of models and use the $n^{\prime}$ labeled models to predict a label for the remaining $n-n^{\prime}$ models. In practice an unlabeled model may not be available when inducing the DKPS or, if $n$ is large, it may be too expensive to induce a DKPS whenever a prediction for a new unlabeled model is required. Out-of-sample techniques (Bengio et al., 2003; Trosset & Priebe, 2008) can be used to meet these imposed constraints at the cost of a slight degradation in representation quality and, hence, inference performance.
@@
Acknowledgements
We would like to thank Avanti Athreya, Tianyi Chen, Ben Johnson, Michael Trosset, Tim Wang, Zekun Wang, and Weiwei Yang for providing helpful comments throughout the development of this manuscript.

### v3

@@ 2.3 An illustrative example – “Was RA Fisher great?"
Figure 1: Left. The 2-d Data Kernel Perspective Space (DKPS) and covariate surface for a collection of 550 models parameterized by fixed augmentations. Right. The performance of the 1-nearest neighbor regressor in DKPS for predicting the probability that an unlabeled model responds “yes" to “Was RA Fisher great?".
@@
The first model-level inference task we study is a toy example where the task is to predict the probability that a model will respond “yes" to the question “Was RA Fisher great?". The question is chosen due to its subjectiveness – there is no correct answer to the question of a person’s greatness – and its duality – Ronald A. Fisher is considered one of the most influential statisticians in history and is considered an advocate of the 20th century Eugenics movement (Bodmer et al., 2021).
@@
We use a 4-bit version of Meta’s LLaMA-2-7B-Chat (Touvron et al., 2023) as a base model and consider a collection of models parameterized by fixed context augmentations. Each augmentation $a_{i}$ contains information related to RA Fisher’s statistical achievements (i.e., $a_{i}$ = “RA Fisher pioneered the principles of the design of experiments") or to his involvement in the eugenics movement (i.e., $a_{i}$ = “RA Fisher’s view on eugenics were primarily based on anecdotes and prejudice.") and is prepended to every query. The covariate corresponding to a given model is calculated by prompting the base model with the appropriately formatted prompt “Give a precise answer to the question based on the context. Don’t be verbose. The answer should be either a yes or a no. $a_{i}$. Was RA Fisher great?" until there are 100 valid responses. We let $y_{i}$ be the average number of “yes"es.
@@
To induce a DKPS for this task, we consider queries sampled from OpenAI’s ChatGPT with the prompt “Provide 100 questions related to RA Fisher.". For a given query $q_{j}$ we prompt the base model with the appropriately formatted prompt “$a_{i}\;q_{j}$" and fix $r=1$. The left figure of Figure 1 is a 3-d figure where the first two dimensions are the DKPS of the $n=550$ ($275$ statistics augmentations, $275$ eugenics augmentations) models induced with $m=100$ queries. Model responses are embedded by averaging the per-token last layer activation of the base model. The third dimension is an interpolated $y_{i}$ surface with a linear kernel (Du Toit, 2008). We emphasize that the dots populating the left figure of Figure 1 are representations of the models with respect to all 100 queries – not just vector representations of a query or a response to a query – and that the $y$-surface corresponds to the average response of each model to a query not included in the 100 queries.
The first dimension of the DKPS is clearly capable of distinguishing between models adorned with “statistics" augmentations and models adorned with “eugenics" augmentations. The shape of the interpolated covariate surface is highly correlated with this feature. A description of the augmentations that parameterize the models and the queries used to induce the DKPS is provided in Appendix B.1.
@@
The right figure of Figure 1 shows the performance of a 1-nearest neighbor (1-NN) regressor in DKPS for a varying number of labeled models and a varying number of queries. The DKPS is induced with $n$ models and the regressor is trained with $n-1$ of the model-level covariates. The mean squared error reported is the error of the 1-NN regressor for predicting the “left out" model’s covariate ($\pm$ 1 S.E.). As Theorems 1 and 2 suggest, the performance of the regressor is dependent on both the number of models and the number of queries: the more models and the more queries the better. The scale of the impact of more models and more queries, however, depends on its counterpart. The amount of models does not have a large impact on predictive performance if the number of queries is small. The amount of queries has a large impact on performance regardless of the number of models.
@@ 3 Experiments
We next consider two experiments with more realistic model-level covariates: predicting whether or not a model has had access to sensitive information and predicting model safety. For all experiments we fix $r=1$ as suggested by the empirical rates of convergence of the perspectives described in Acharyya et al. (2024). We use the MDS implementation from Graspologic (Chung et al., 2019) throughout. We use the profile likelihood of the singular values of $D$ to determine the dimensionality of the DKPS (Zhu and Ghodsi, 2006) and note that this may be larger than two. We show only the first two dimensions for visualization purposes.
@@ 3.1 Has a model seen sensitive information?
Figure 2: Left. The 2-d data kernel perspective space (DKPS) of 50 fine-tuned models – 25 with “sensitive" data in the fine-tuning data mixture (red), 25 with none (black) – induced by an evaluation set containing 10 prompts relevant to the sensitive data. For models trained on sensitive data, color intensity correlates with amount of sensitive data in the training mixture. Center. The 2-d DKPS of the models induced by a set of 10 prompts “orthogonal" to the difference between models with sensitive data in their fine-tuning data mixture and models with no sensitive data in their fine-tuning data mixture. Right. Classification performance as a function of number of labeled models and size of evaluation set for both sensitive and orthogonal evaluation sets.
Modern language models are trained with trillions of tokens of text (Touvron et al., 2023). For proprietary models such as OpenAI’s GPT series or Anthropic’s Claude series, the exact sources of the training mixtures are unknown and – given the models’ propensities to produce content that is strikingly similar to copyrighted content (Henderson et al., 2023) – its curation and use is ethically questionable (Lemley and Casey, 2020). Further, the training mixture of some models may include sensitive informative such as personal information, trade secrets, or government-classified information that should never be presented to the end user. Developing classifiers to identify models that are either trained on sensitive or copyrighted information or are likely to produce sensitive or copyrighted information is thus paramount to uphold the rights of the stakeholders of the original content.
To investigate the utility of DKPS for this purpose, we again use a 4-bit version of LLaMA-2-7B-Chat as a base model and train 50 different LoRA adapters with different subsets of the Yahoo! Answers (YA) dataset (Zhang et al., 2015). The YA dataset consists of data from 10 topics. We consider data from the topic “Politics & Government" to be “sensitive" information, data from the topics “Society & Culture", “Science & Mathematics", “Health", “Education & Reference", “Computers & Internet", and “Sports" to be “not-sensitive", and data from the remaining topics to be “orthogonal". We trained each of the 50 adapters with 500 question-answer pairs for 3 epochs with a learning rate of $5\times 10^{-5}$ and a batch size of 8. Each adapter is rank 8, has a scaling factor of 32, targets all attention layers, does not have bias terms, and has a dropout probability of 0.05 when training. For 25 of the adapters, the adapter training mixture consisted wholly of randomly selected not-sensitive data. For the remaining 25 adapters, the adapter training mixture consisted of $p_{i}$ randomly selected sensitive data and $500-p_{i}$ randomly selected not-sensitive data. We let $y_{i}$ be the indicator of whether or not the adapter’s training mixture contained any sensitive data.
@@
models and a randomly selected query set. We report the expected risk of the classifier trained on a random subset of the models for the remaining models. We observe similar phenomena to the “Was RA Fisher great?" experiment in that both the amount of models and the amount of queries impact performance. For a fixed $m$, the expected risk curves also highlight the observed difference in separability of the models in the two DKPS, with the expected risk with sensitive queries being significantly lower than the expected risk with orthogonal queries. Indeed, the expected risk with $m=10$ sensitive queries is similar to the expected risk with $m=50$ orthogonal queries.
@@
Throughout this experiment we use the term “orthogonal" for queries from topics that are not used when fine-tuning the models under study. The term is chosen because, naïvely, queries from these topics should not elicit different responses from models trained on sensitive data and models trained on not-sensitive data. In reality, the sensitive data and the not-sensitive data have different underlying token distributions and this difference will cause systematic differences in model responses after fine-tuning even for topics that are irrelevant a priori. Further, it is likely that the documents in the “orthogonal" topics share some content commonalities with documents in the sensitive and not-sensitive topics. We see this phenomenon in the classification results in the right figure of Figure 2 where the linear classifier is able to perform better than chance with enough “orthogonal" queries.
@@ 3.2 How safe is a model?
Figure 4: Left. A graph where each node is a model and an edge between two models exists if model $i$ is fine-tuned from model $i^{\prime}$ or if model $i^{\prime}$s weights were used in a model-merge that resulted in model $i$, etc. Right, top. The two-dimensional data kernel perspective spaces (DKPS) corresponding to the toxicity and bias prediction tasks. Dot size is proportional to model toxicity or bias. Right, bottom. Relative performance of three regression techniques.
Model safety is one of the main factors when developing a language model in production. An unsafe model is prone to propagating harmful stereotypes (Ferrara, 2024), using toxic language (Wen et al., 2023), and misunderstanding the user’s intent (Ji et al., 2023) – all of which can adversely affect the user and their experience. Hence, developing techniques
to understand how unsafe a model is an important aspect of the model-production pipeline. As with predicting if a model has had access to sensitive information above, we investigate using DKPS to predict model safety through the lens of model toxicity and model bias.
@@
For each model in the collection we consider two covariates: model toxicity and model bias. To determine a model’s toxicity, we prompt each model with a collection of queries from the Real Toxicity Prompts (RTP) (Gehman et al., 2020) dataset and subsequently evaluate the toxicity of each response with the neural model roberta-hate-speech-dynabench-r4 (Vidgen et al., 2021). The model-level toxicity is simply the average response toxicity. An analogous process is used to define a model’s bias with the dataset Bias in Open-ended Language Generation Dataset (BOLD) (Dhamala et al., 2021) and the regard model (Sheng et al., 2019).
@@
We quantify this observation by evaluating regressors for predicting the model-level toxicity and bias of unscored models. We consider three different regressors for these tasks. The first is a constant equal to the average covariate of the scored models (i.e., the “global mean"). The second uses the average covariate of models who share an edge with the unscored model (i.e., “1-NN (graph)"). The third uses the covariate of the nearest neighbor in DKPS (i.e., “1-NN (DKPS)"). For 1-NN (DKPS) we consider varying amounts of randomly sampled queries from RTP and BOLD to induce the DKPS. We use the global mean as a standard and report the relative absolute error of the three methods in the bottom right of Figure 4. The reported performance of the 1-NN (DKPS) regressor is the average of the smaller of $200$ and $2000/m$ random samples of $m$ queries over all possible leave-one-model-out scenarios. Notably, given enough queries local predictions in DKPS outperform local predictions in the graph space and predictions using the global mean.
In addition to reporting the relative performance, we report the time it takes to use the DKPS machinery to predict Mistralv1.0’s (Jiang et al., 2023) toxicity and bias relative to using the evaluation model through HuggingFace’s API (Wolf et al., 2019). The time we report for the HuggingFace API is the time it takes to calculate the “ground truth" used to calculate the performance of the regressors above and is the total time required for Mistralv1.0 to produce responses and for the evaluation model to produce scores for 2000 queries. The time we report for 1-NN (DKPS) includes the time it takes for Mistralv1.0 to produce $m$ responses, the time it takes nomic-embed-v1.5 to embed the responses, the time it takes to induce the DKPS, and the time it takes to train and use the nearest neighbor regressor. It does not include the time it takes to produce and evaluate the responses of the other models in the collection. The relative efficiency of DKPS, as seen in Figure 5, is approximately $1/m$. This relationship will hold for all model-level inference tasks where the covariate is proportional to the sum of a function of individual responses.
@@ 4 Discussion
We have demonstrated – both theoretically and empirically – that embedding-based representations of generative models can be used for various model-level inference tasks. While the results we presented show the potential of our approach, there are choices throughout the collection-of-models-to-covariate-prediction pipeline that can affect performance and implementation practicality. We outline these choices in the Limitations section.
5 Limitations
@@
As mentioned in Section 2.2.1, one major decision is the collection of queries $q_{1},,\ldots,q_{m}$ used to induce the perspective space. In particular, the representations of the models that the query set induces may or may not be relevant to a particular inference task. We demonstrate this phenomenon in Section 3.1 where the queries from the “sensitive-only" query distribution can induce representations of similar discriminative ability as queries from the “orthogonal-only" query distribution with $1/5$ of the queries. We expect a similar but less dramatic effect within the distributions of “sensitive-only" queries and anticipate that curating an “optimal" set of queries for a given $g$ and for a fixed number of queries will soon be a highly active research area.
@@
Another decision is the distance function used to define $D$. The MDS of the distance matrix studied herein (Eq. (1)) produces representations of the models that are consistent for objects that capture the true mean discrepancy geometry of the model responses. For model-level covariates that cannot naturally be described as a function of the mean discrepancy geometry, we do not expect that information-theoretic optimal performance is possible for any $n,m,$ and $r$, even where the queries are from an “optimal" query distribution. Instead, for representations of the models that can be used generally without information loss, it may be necessary to replace the Frobenius norm of the differences of the average embedded response with a more expressive distance such as one defined directly on the cumulative distributions of responses or with a task-specific distance (Helm et al., 2020). Related theoretical work (Tang et al., 2013) suggests that our results will hold for more expressive distances. We do not expect the naïve replacement of the Frobenius norm with a more expressive distance function to be universally better for model-level inference tasks in practice as there are likely computational and query set quality trade-offs to consider. As with the active curation of an optimal query set, we expect these trade-offs to be important future research topics.
@@
In the experiments above we have access to $n$ models and $n^{\prime}<n$ corresponding model-level covariates. We induce a DKPS with the entire collection of models and use the $n^{\prime}$ labeled models to predict a label for the remaining $n-n^{\prime}$ models. In practice an unlabeled model may not be available when inducing the DKPS or, if $n$ is large, it may be too expensive to induce a DKPS whenever a prediction for a new unlabeled model is required. Out-of-sample techniques (Bengio et al., 2003; Trosset and Priebe, 2008) can be used to meet these imposed constraints at the cost of a slight degradation in representation quality and, hence, inference performance.